OpenStudy (anonymous):

Consider the functions: f=1/(2x-4) and g=2/(x-3); Calculate f/g and find its Domain(f/g)

OpenStudy (anonymous):

\[\frac{1}{2x-4}\div \frac{2}{x-3}=\frac{1}{2x-4}\times \frac{x-3}{2}\] is a start

OpenStudy (anonymous):

as for the domain, you have to exclude the solutions to \(2x-4=0\) as well as \(x-3=0\)

OpenStudy (anonymous):

yes that part i get. so the answer is (x-3)/4x-8

OpenStudy (anonymous):

why excluding x-3=0 though, that's where i'm lost

OpenStudy (anonymous):

x=2 would make the denominator vanish but why can't x be 3